OpenStudy (anonymous):

what is the simplest form of the quptient 3 sqrt 162 / 3 sqrt 2

OpenStudy (anonymous):

do you know how to start this problem?

OpenStudy (anonymous):

no im having problems with these

OpenStudy (anonymous):

you have to find factors of 162 that are perfect squares. whats the biggest perfect square that is a factor of 162?

OpenStudy (anonymous):

do you mean 54

OpenStudy (anonymous):

54 isnt a perfect square perfect squares = 1,4,9,16, etc

OpenStudy (anonymous):

9

OpenStudy (anonymous):

i was dividing by 3

OpenStudy (anonymous):

you can factor sqrt(162) into sqrt(81) times sqrt(2). sqrt(81)=9, the sqrt(2) on the top and bottom cancels, then simplify from there

OpenStudy (anonymous):

so how would you write this problem out

OpenStudy (anonymous):

if you could show me

OpenStudy (anonymous):

\[\frac{ 3\sqrt{162} }{ 3\sqrt{2} }\] the 3 on the top and bottom cancel \[\sqrt{162}=\sqrt{81}\sqrt{2}=9\sqrt{2}\] the \[\sqrt{2}\] on the top and bottom cancel, and the answer is 9

OpenStudy (anonymous):

so actually you ignore the 3 and divide 162 by 2 to get 81 then find the sqrt of 81 which is 9

OpenStudy (anonymous):

yes. then you divide the 3 on top by the 3 on bottom and the sqrt(2) on top by the sqrt(2) on the bottom, and youre left with 9

OpenStudy (anonymous):

okay kinda sounds easy enough thank you

OpenStudy (anonymous):

welcome:)